OpenStudy (anonymous):

The function f is continuous on the interval [3, 13] with selected values of x and f(x) given in the table below. Use the data in the table to approximate f ′(3.5).

OpenStudy (anonymous):

x 3 4 7 10 13 f(x) 2 8 10 12 22

OpenStudy (anonymous):

f(3.5) is in the middle of f(3) and f(4), so you do a simple average/mean equation. (2+8)/2=10/2=5

OpenStudy (anonymous):

Alright, I got that, but I am unsure on how to get the derivative of it without an equation of some sort.

OpenStudy (anonymous):

@ijakez

OpenStudy (anonymous):

OH, didn't see the f prime! My mistake! It would be: since f(3.5) is in the middle of f(4) and f(3), you do: [f(4)-f(3)]/(4-3) which is (8-2)/1=6

OpenStudy (anonymous):

How does that yield the f prime?

OpenStudy (anonymous):

Wait, is that supposed to be change in y with change in x.

OpenStudy (anonymous):

I don't think that would be it.

OpenStudy (anonymous):

Wait, you were right, thanks! Do you mind answering another question?

OpenStudy (anonymous):

f(x) and g(x) are a differentiable function for all reals and h(x) = g[f(3x)]. The table below gives selected values for f(x), g(x), f '(x), and g '(x). Find the value of h '(1).

OpenStudy (anonymous):

x 1 2 3 4 5 6 f(x) 0 3 2 1 2 0 g(x) 1 3 2 6 5 0 f '(x) 3 2 1 4 0 2 g '(x) 1 5 4 3 2 0

OpenStudy (anonymous):

@ijakez

OpenStudy (anonymous):

I'm back! Give me a minute

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I solved it and got 15. I am not sure if I am correct though.

OpenStudy (anonymous):

My working was g'(f(3x)) x f'(3x) x 3 g'(2) x f'(3) x 3 5 x 1 x 3 15

OpenStudy (anonymous):

How did you get g'(f(2))

OpenStudy (anonymous):

I think I'm right, I'm going to try my answer.

OpenStudy (anonymous):

Tell me how it goes!

OpenStudy (anonymous):

I was right! thanks anyway though, your time was greatly appreciated!!!!

OpenStudy (anonymous):

Sure, and great job!

OpenStudy (anonymous):

I'd give you another medal, but I can't :(

OpenStudy (anonymous):

All good! I'm in calc too, so I need as much practice as I can get!

OpenStudy (anonymous):

Yeah, the AP exam is in a month!